# Vision-and-Language Training Helps Deploy Taxonomic Knowledge but Does Not Fundamentally Alter It

### v1

@@ 1 Introduction
Many standard VLMs [32, 29, i.a.] often build on top of a pretrained language model (LM) by adding visual conditioning to its next token prediction task, often also updating the parameters of the language model. Analyses of VLM capabilities often focus on the multimodal tasks this additional modality enables. But (how) does this vision-and-language (VL) training change the linguistic capacity of the model? Answering this question requires comparing VL-tuned LMs to their original LM counterparts. Empirical evidence in this literature is rather sparse, often comparing such “VLM-LM minimal pairs” on general benchmarks such as MMLU [16] and GLUE [63]. In this paper, we consider a more targeted investigation (like [71]) of VLM-LM pairs in a particular domain: lexical-conceptual knowledge, specifically its taxonomic organization (e.g., a cat is an animal). Evaluation of taxonomic knowledge has been of continued interest within the Natural Language Processing [14, 31, 42, 44] and Computer Vision communities [2, 61, 47]—however, to the best of our knowledge no work so far has compared minimally differing VLM-LM pairs in terms of how well they can reason taxonomically.
@@ 2 Related Work
There are two main strands of empirical work measuring the influence of the additional visual modality on models’ linguistic behavior and representations. The first line of work compares VLM and LM performance on downstream text-only benchmarks. The results are mixed: for instance, FLAVA [59] noted around 8% point gains over the base masked language model on GLUE-style NLP tasks (although the evaluation setting involved finetuning). On the other hand, Molmo has been reported to be outperformed by its base LM, Qwen, on text-only benchmarks like MMLU [8]. Generally, more evidence exists in favor of multimodal training hurting text-only task performance [20, 36] and this observation has been used to argue for freezing the language part of the model during multimodal training [12]. The second line of work conducts more targeted comparisons of VLMs and LMs, examining whether additional vision training leads to differences in representations of syntactic categories [64] and performance on tasks that require more “grounding” [71], but the findings overall have indicated no substantial differences.
@@ 3.1 Dataset design
We created a QA dataset that requires taxonomic understanding based on GQA [18], named TaxonomiGQA. A datapoint in GQA consists of an image of a scene, a question about this scene, and metadata that includes a scene graph of the objects, their attributes, and relations between the objects. We applied a three-step modification (each step illustrated in Figure 1) to create a text-only dataset, since our goal is to systematically compare the LM and VLMs’ taxonomic competence. (1) Convert the scene graph into a purely textual description of the scene programmatically using hand-crafted templates; (2) For each question that contains a word that corresponds to a node in our reference taxonomy, substitute the word to its hypernym; (3) For each substitution, create four negative samples following Misra et al. [42], by substituting the target word with a word that is not in its hypernymy chain. Below we describe the resources, filtering, and sampling details used to create TaxonomiGQA.
@@
To construct the reference taxonomy, we first extracted all unique noun lemmas ($N=1216$) that appeared in the GQA questions, and annotated their senses in the WordNet taxonomy, these serve as the leaf nodes of our taxonomy. Next, for each noun, we extracted its hypernym chain (e.g., dog < canine < mammal < vertebrate < animal) from WordNet, rejecting hypernyms that were too abstract (determined manually), e.g., entity, material, conveyance. 315 concepts were removed as a result, many of which often had abstract entities in their hypernym chains or had non-ideal WordNet categorization (e.g., bubble as a member of ball), leaving us with 901 unique chains.
@@
We applied a multi-stage filtering process to the validation split of GQA (10,696 images/scenes and 488,293 questions) to obtain our base questions. We first applied scene-level filtering by excluding scenes containing more than 20 annotated objects or any repeated object labels to avoid ambiguity in referring expressions in text. For each remaining scene, we applied question-level filtering to retain questions that refer to a single object (excluding any that mention multiple objects) and whose hypernyms do not overlap with those of any other object in the scene. Next, we balanced the dataset by randomly sampling 40 questions per scene in proportion to each scene’s question type ratios. We further filtered the questions by answer type and restricted the dataset to yes/no questions to facilitate the substitution step. This reduced our taxonomy to 314 unique chains. In the base questions remaining after filtering, we substituted each target concept with each of its hypernyms in its hypernym chain to obtain the substituted questions. Then, we created negative sample questions by substituting the target concepts with concepts that are not in their hypernym chain, discarding question types where this substitution was not possible due to the introduction of presupposition failure (e.g., questions such as Is the color of the dog brown? when there is no dog in the image). This ended up eliminating more hypernym chains (which were only present in the discarded question types), leaving us with 126 final chains. More details about this negative sampling pipeline is given in Appendix C.
@@
The final dataset contains 1,342 unique images/scenes, 29,604 positive sample instances (9,334 targeting leaf node concepts, 20,270 targeting hypernym-substitutions), and 4 negative samples for each positive sample, amounting to 148,020 total instances. There are 276 hyponym-hypernym pairs, 126 unique hypernym chains, 88 unique hypernyms, and 24 top-level categories.
@@ 3.2 Metrics
[…]
@@
proposed by Wu et al. [66, originally named “Hierarchical Consistence Accuracy”] measures a stricter form of accuracy relative to the previous ones, as the proportion of time the model correctly answers the original unsubstituted instance and all of its corresponding hypernym-substituted instances. Using our notation, this is measured as:
@@ 3.4 Results
The results are shown in Figure 2: points above the diagonal denote model pairs where the VLM outperforms the LM counterpart, and points below denote model pairs where the LM outperforms the VLM. We observe a consistent trend (with a single exception of Vicuna vs. Llava-1.5) where the VLMs outperform their LM counterpart, even though the presentation of the task was purely linguistic. [^1] For the curious: see Appendix D for how the VL models perform on our text-only QA vs. VQA. We rule out the possibility that the VLMs are performing better due to having been trained on GQA directly by running a control experiment where we give only the question (without the scene description) to the VLMs—if the VLMs have already seen the original GQA data, they may have learned the question-label association. The results (Appendix D) show that most VLMs do not perform substantially above chance. Since the text descriptions are newly introduced in TaxonomiGQA, we can safely rule out the hypothesis that VLMs’ improvements are due to having been trained on GQA.
Accepting the trend of VLMs outperforming LMs on TaxonomiGQA as a genuine improvement, we conduct analyses that aim to explain this result more. When we are not analyzing all model pairs, we focus on Qwen 2.5-Instruct vs. Qwen 2.5-VL-Instruct, since the performance gap between VLM and LM was the most salient with this pair especially for the stricter metrics (conditional and HC).
Figure 2: Performance of VLM-LM model pairs on TaxonomiGQA (Section 3) and TAXOMPS (Section 4.1). Points above the line indicate that VLM outperforms LM.
@@ 4 H1: VLMs’ taxonomic knowledge aligns better with reference taxonomy
One possible reason that VLMs are performing better on TaxonomiGQA may be due to the difference in their underlying (task-agnostic) taxonomic knowledge, and in particular, in a way that better aligns with the reference taxonomy used to create the hypernym-substituted questions. We test this hypothesis about taxonomic knowledge difference in three different ways: (1) through a QA task that directly elicits taxonomic judgments; (2) through an analysis of the hierarchical organization of concepts in the models’ representation space; and (3) through similarity analysis on the embeddings.
@@ 4.1 Directly eliciting taxonomic judgments through Taxonomic Minimal Pairs (TAXOMPS)
Since TaxonomiGQA presupposes taxonomic knowledge rather than eliciting it directly, we first check whether VLMs and LMs differ in their ability to directly answer questions about taxonomic relations. To this end, we introduce TAXOMPS (Taxonomic Minimal Pairs), a dataset which consists of questions of the form “Is it true that a $\bm{C_{1}}$ is a $\bm{C_{2}}$?” where $\bm{C_{1}}$ (cat) and $\bm{C_{2}}$ (feline) are concepts that are in a hypernymy relation, and negative samples where $\bm{C_{2}}$ is replaced by a concept that is not the hypernym of $\bm{C_{1}}$ (vehicle), following Misra et al. [42]. As in TaxonomiGQA, we use 4 negative samples. We construct TAXOMPS directly from the final taxonomy used in our TaxonomiGQA analysis—i.e., 276 total hyponym-hypernym pairs, each coupled with 4 negative samples, yielding 1380 questions. The metric follows Section 3 (specifically, the Overall Accuracy measure, because there is no conditional analog). That is, an instance is considered correct iff. the model answers questions with the hyponym-hypernym pairs (Is it true that a cat is an animal?) with a Yes while answering No to the negative sample questions (Is it true that a cat is a vehicle/fruit/tool/vegetable?).
Figure 2 shows the results. With the exception of Llama-3.1 vs. MLlama 3.2, most VLM-LM pairs perform quite similarly (and well) on TAXOMPS. This suggests that additional VL training does not in general alter the basic taxonomic membership judgments of a language model.
@@ 4.2 Lexical representations of taxonomic knowledge
Can the alignment with reference taxonomy be observed representationally, although not by direct elicitation? We test whether the lexical representations in VLMs align better with the reference taxonomy than LMs via their hierarchical organization and hypernym-hyponym embedding similarity.
@@ 4.2.1 Hierarchical taxonomic structure
Method
We applied Park et al.’s method to transform the unembedding space in our models to a space where the inner product between two concepts’ vectors is sensitive to the hierarchical relation between them.Then, we compare VLM-LM pairs in terms of their pairwise cosine similarities between concepts in their unembeddings’ causal inner product space (as established in [52]). In addition, we use the large WordNet hierarchy (a superset of our taxonomy) originally used by [51] to compare the pairwise similarities of concepts in VLM and LM to that of the pairwise path-similarities between concepts in WordNet. These comparisons are done using Representational Similarity Analysis [24], which computes the Spearman’s correlation between two matrices’ (flattened) upper triangular matrices, and take these values as the representational similarity between the three spaces: VLM, LM, WordNet. Greater RSA value between two spaces indicates greater similarity between. To account for potential variance, we sampled 100 subsets (of size 100 $\times$ 100 each) from the larger pairwise matrices and report the mean and standard deviation of the RSA correlations across all subsets.
Results
This analysis (Table 1, left) shows that the hierarchical organization of concepts (as defined by [51]) is mostly shared between the VLM and LM, indicated by the consistently similar RSA scores when comparing VLMs and LMs to WordNet, as well as the high similarity between the VLM and LM when directly compared (all RSA scores $\geq$ 0.95). Interestingly, the Qwen 2.5 and Molmo pairs, the two model pairs that showed the most salient advantage of VLMs in Figure 2 had the lowest VLM-LM RSA scores: 0.95 and 0.96, respectively. However these values are still very high in terms of raw correlation, suggesting that they are still fundamentally similar. The pairwise similarities for VLMs, LMs, and WordNet can be visually inspected in Figure 5 in Appendix E.
Table 1: Left: RSA comparisons of hierarchically sensitive pairwise similarities [51] in the unembedding spaces of VLM-LM pairs, and pairwise path-similarities from the WordNet (wn) Noun Hierarchy. Subscripts show standard deviation (hidden if under 0.01). Right: Differences ($\Delta$) in cosine similarities between positive concept pairs (i.e., in a hypernymy relationship) to those in negative samples from the taxonomy in TaxonomiGQA, computed using VLM and LM static-embedding layers
@@ 4.2.2 Embedding similarities of taxonomic relations
Another way in which taxonomic relations can be encoded is via vector similarity—we test the hypothesis that the lexical embeddings (i.e., uncontextualized representations) corresponding to concepts in our reference taxonomy are more similar to embeddings of their hyponyms, compared to embeddings of non-hyponyms in our taxonomy. We compute the similarity between a target concept and four randomly sampled non-hyponym concepts (same as in Section 3.1). Then, we compute the difference between target-hyponym similarity and the average similarity between the target and the negative samples. We test whether this difference is greater in VLMs than LMs, which would mean that VLM embeddings encode hypernym-hyponym relations more similarly than non-hypernym-hyponym relations. This hypothesis is not borne out: Table 1 (right) shows that this holds for no VLM-LM pairs (the significant effect in Qwen2.5-I vs. Qwen2.5-VL-I is in the opposite direction).

### v2

@@ 1 Introduction
Many standard VLMs [33, 30, i.a.] often build on top of a pretrained language model (LM) by adding visual conditioning to its next token prediction task, often also updating the parameters of the language model. Analyses of VLM capabilities often focus on the multimodal tasks this additional modality enables. But (how) does this vision-and-language (VL) training change the linguistic capacity of the model? Answering this question requires comparing VL-tuned LMs to their original LM counterparts. Empirical evidence in this literature is rather sparse, often comparing such “VLM-LM minimal pairs” on general benchmarks such as MMLU [17] and GLUE [65]. In this paper, we consider a more targeted investigation (like [73]) of VLM-LM pairs in a particular domain: lexical-conceptual knowledge, specifically its taxonomic organization (e.g., a cat is an animal). Evaluation of taxonomic knowledge has been of continued interest within the Natural Language Processing [14, 32, 43, 45] and Computer Vision communities [2, 62, 48]—however, to the best of our knowledge no work so far has compared minimally differing VLM-LM pairs in terms of how well they can reason taxonomically.
@@ 2 Related Work
There are two main strands of empirical work measuring the influence of the additional visual modality on models’ linguistic behavior and representations. The first line of work compares VLM and LM performance on downstream text-only benchmarks. The results are mixed: for instance, FLAVA [60] noted around 8% point gains over the base masked language model on GLUE-style NLP tasks (although the evaluation setting involved finetuning). On the other hand, Molmo has been reported to be outperformed by its base LM, Qwen, on text-only benchmarks like MMLU [8]. Generally, more evidence exists in favor of multimodal training hurting text-only task performance [21, 37] and this observation has been used to argue for freezing the language part of the model during multimodal training [12]. The second line of work conducts more targeted comparisons of VLMs and LMs, examining whether additional vision training leads to differences in representations of syntactic categories [66] and performance on tasks that require more “grounding” [73], but the findings overall have indicated no substantial differences. We contribute to this line of work by showcasing a context where there is a non-trivial difference brought about as a result of VL training. In particular, we show that while VL training does not fundamentally alter task-agnostic representations of taxonomic knowledge in LMs (in line with prior work), it does improve the deployment of this knowledge in the context of a question-answering task.
@@ 3.1 Dataset design
We created a QA dataset that requires taxonomic understanding based on GQA [19], named TaxonomiGQA. A datapoint in GQA consists of an image of a scene, a question about this scene, and metadata that includes a scene graph of the objects, their attributes, and relations between the objects. We applied a three-step modification (each step illustrated in Figure˜1) to create a text-only dataset, since our goal is to systematically compare the LM and VLMs’ taxonomic competence. (1) Convert the scene graph into a purely textual description of the scene programmatically using hand-crafted templates; (2) For each question that contains a word that corresponds to a node in our reference taxonomy, substitute the word to its hypernym; (3) For each substitution, create four negative samples following Misra et al. [43], by substituting the target word with a word that is not in its hypernymy chain. Below we describe the resources, filtering, and sampling details used to create TaxonomiGQA.
@@
To construct the reference taxonomy, we first extracted all unique noun lemmas ($N=1216$) that appeared in the GQA questions, and annotated their senses in the WordNet taxonomy, these serve as the leaf nodes of our taxonomy. Next, for each noun, we extracted its hypernym chain (e.g., dog < canine < mammal < vertebrate < animal) from WordNet, rejecting hypernyms that were too abstract (determined manually), e.g., entity, material, conveyance. 315 concepts were removed as a result, many of which often had abstract entities in their hypernym chains or had non-ideal WordNet categorization (e.g., bubble as a member of ball), leaving us with 901 unique chains. (See Appendix˜C for more details.)
@@
We applied a multi-stage filtering process to the validation split of GQA (10,696 images/scenes and 488,293 questions) to obtain our base questions. We first applied scene-level filtering by excluding scenes containing more than 20 annotated objects or any repeated object labels to avoid ambiguity in referring expressions in text. For each remaining scene, we applied question-level filtering to retain questions that refer to a single object (excluding any that mention multiple objects) and whose hypernyms do not overlap with those of any other object in the scene. Next, we balanced the dataset by randomly sampling 40 questions per scene in proportion to each scene’s question type ratios. We further filtered the questions by answer type and restricted the dataset to yes/no questions to facilitate the substitution step. This reduced our taxonomy to 314 unique chains. In the base questions remaining after filtering, we substituted each target concept with each of its hypernyms in its hypernym chain to obtain the substituted questions. Then, we created negative sample questions by substituting the target concepts with concepts that are not in their hypernym chain, discarding question types where this substitution was not possible due to the introduction of presupposition failure (e.g., questions such as Is the color of the dog brown? when there is no dog in the image). This ended up eliminating more hypernym chains (which were only present in the discarded question types), leaving us with 126 final chains. More details about this negative sampling pipeline is given in Appendix D.
@@
The final dataset contains 1,342 unique images/scenes, 29,604 positive sample instances (9,334 targeting leaf node concepts, 20,270 targeting hypernym-substitutions), and 4 negative samples for each positive sample, amounting to 148,020 total instances. There are 276 hyponym-hypernym pairs, 126 unique hypernym chains, 88 unique hypernyms, and 24 top-level categories (e.g., animal, vehicle, etc.).
@@ 3.2 Metrics
As preliminaries, each instance, $X_{i}=(q,q^{n}_{1...4})$ in TaxonomiGQA consists of a positive sample question, $q$, about some leaf-level category (target concept), coupled with a set of 4 negative sample questions, $\{q^{n}_{1...4}\}$, where the target concept in the original question is now replaced by a negative-sample concept, as described in Section˜3. Next, for each instance, we have a set of $k_{i}$ hypernym-substituted instances, $\{X^{s,i}_{1},\dots,X_{k_{i}}^{s}\}$, where each item $X^{s}_{j}$ is an instance but with the original category substituted with a category in its hypernym chain, along with their own 4 negative samples. Finally, we use a function, $\texttt{correct(.)}\rightarrow\{0,1\}$, which accepts an instance $X$, and returns $1$ iff. the model correctly answers the positive sample question and all four negative sample questions, and 0 otherwise. Using these preliminaries, we propose the following metrics:
@@
proposed by Wu et al. [68, originally named “Hierarchical Consistence Accuracy”] measures a stricter form of accuracy relative to the previous ones, as the proportion of time the model correctly answers the original unsubstituted instance and all of its corresponding hypernym-substituted instances. Using our notation, this is measured as:
@@ 3.4 Results
The results are shown in Figure 2: points above the diagonal denote model pairs where the VLM outperforms the LM counterpart, and points below denote model pairs where the LM outperforms the VLM. We observe a consistent trend (with a single exception of Vicuna vs. Llava-1.5) where the VLMs outperform their LM counterpart, even though the presentation of the task was purely linguistic. [^1] For the curious: see Appendix E for how the VL models perform on our text-only QA vs. VQA. We rule out the possibility that the VLMs are performing better due to having been trained on GQA directly by running a control experiment where we give only the question (without the scene description) to the VLMs—if VLMs have encountered the original GQA instances during training, they may have learned question-label associations used in our dataset. Our results (Appendix˜E) show that most VLMs do not perform substantially above chance. Since the text descriptions are newly introduced in TaxonomiGQA, we can safely rule out the hypothesis that VLMs’ improvements are due to having been trained on GQA. Accepting the trend of VLMs outperforming LMs on TaxonomiGQA as a genuine improvement, we conduct analyses that aim to explain this result more in the subsequent sections. When we are not analyzing all model pairs, we focus on Qwen 2.5-Instruct vs. Qwen 2.5-VL-Instruct, since the performance gap between VLM and LM was the most salient with this pair, especially on stricter metrics (conditional accuracy and HC).
Generalizability of our finding
We conducted a supplementary experiment using the dataset from Rodriguez et al. [56]—another case where task contexts require the deployment of taxonomic knowledge—to verify the generalizability of our finding outside of TaxonomiGQA. Here, LMs were evaluated on their projection of novel properties (e.g., is daxable, has feps, etc.) from hypernyms to their hyponyms. Results on this dataset (shown in Table˜5 in Appendix˜F) are qualitatively in line with those on TaxonomiGQA: VLMs were substantially better than their LM counterparts in 5 out of 7 VLM-LM pairs. Furthermore, concurrent work by Tan et al. [63] that also investigates taxonomic understanding in minimally differing LMs and VLMs corroborates our finding. In their results, VLMs (with the exception of LlaVa-OneVision, and to a milder extent, InternVL-8B) showed improved performance on 4 out of 5 taxonomic understanding benchmarks.
Figure 2: Performance of VLM-LM model pairs on TaxonomiGQA (Section˜3) and TAXOMPS (Section˜4.1). Points above the line indicate that VLM outperforms LM.
@@ 4 H1: VLMs’ taxonomic knowledge aligns better with reference taxonomy
One possible reason that VLMs are performing better on TaxonomiGQA could be due to the difference in their underlying (task-agnostic) taxonomic knowledge, and in particular, in a way that better aligns with the reference taxonomy used to create the hypernym-substituted questions. We test this hypothesis about taxonomic knowledge difference in three different ways: (1) through a QA task that directly elicits taxonomic judgments; (2) through an analysis of the hierarchical organization of concepts in the models’ representation space; and (3) through similarity analysis on the embeddings.
@@ 4.1 Directly eliciting taxonomic judgments through Taxonomic Minimal Pairs (TAXOMPS)
Since TaxonomiGQA presupposes taxonomic knowledge rather than eliciting it directly, we first checked whether VLMs and LMs differed in their ability to directly answer questions about taxonomic relations. To this end, we introduce TAXOMPS (Taxonomic Minimal Pairs), a dataset which consists of questions of the form “Is it true that a $\bm{C_{1}}$ is a $\bm{C_{2}}$?” where $\bm{C_{1}}$ (cat) and $\bm{C_{2}}$ (feline) are concepts that are in a hypernymy relation, and negative samples where $\bm{C_{2}}$ is replaced by a concept that is not the hypernym of $\bm{C_{1}}$ (vehicle), following Misra et al. [43]. We constructed TAXOMPS directly from the final taxonomy used in our TaxonomiGQA analysis—i.e., 276 total hyponym-hypernym pairs, each coupled with 4 negative samples (same as in TaxonomiGQA), yielding 1380 questions. We use Overall Accuracy as our performance measure (since there is no conditional analog), following Section˜3. That is, an instance is considered correct iff. the model answers questions with the hyponym-hypernym pairs (Is it true that a cat is an animal?) with a Yes while answering No to the negative sample questions (Is it true that a cat is a vehicle/fruit/tool/vegetable?).
Figure˜2 shows our results. With the exception of Llama-3.1 vs. MLlama 3.2, most VLM-LM pairs perform quite similarly (and well) on TAXOMPS. This suggests that additional VL training does not in general alter the basic taxonomic membership judgments of a language model.
@@ 4.2 Lexical representations of taxonomic knowledge
Can the alignment with reference taxonomy be observed representationally, although not by direct elicitation? We tested whether the lexical representations in VLMs align better with the reference taxonomy than LMs via their hierarchical organization and hypernym-hyponym embedding similarity.
@@ 4.2.1 Hierarchical taxonomic structure
We applied Park et al. [52]’s method to transform the unembedding space in our models to a space where the inner product between two concepts’ vectors is sensitive to the hierarchical relation between them. Then, we compared VLM-LM pairs in terms of their pairwise cosine similarities between concepts in their unembeddings’ causal inner product space (as established in [53]). In addition, we used the large WordNet hierarchy (a superset of our taxonomy) originally used by [52] to compare the pairwise similarities of concepts in VLM and LM to that of the pairwise path-similarities between concepts in WordNet. We conducted these comparisons using Representational Similarity Analysis [25], which computes the Spearman’s correlation between two matrices’ (flattened) upper triangular matrices, treating it as the representational similarity between the two spaces. We conducted RSA between three representational spaces: VLM, LM, WordNet. Greater RSA value between two spaces indicates greater similarity between. To account for potential variance, we sampled 100 subsets (of size 100 $\times$ 100 each) from the full pairwise matrices and report the mean and standard deviation of the RSA correlations across all subsets.
This analysis (Table˜1, left) shows that the hierarchical organization of concepts (as defined by [52]) is mostly shared between the VLM and LM, indicated by the consistently similar RSA scores when comparing VLMs and LMs to WordNet, as well as the high similarity between the VLM and LM when directly compared (all RSA scores $\geq$ 0.95). Interestingly, the Qwen 2.5 and Molmo pairs, the two model pairs that showed the most salient advantage of VLMs in Figure˜2 had the lowest VLM-LM RSA scores: 0.95 and 0.96, respectively. However these values are still very high in terms of raw correlation, suggesting that they are still fundamentally similar. The pairwise similarities for VLMs, LMs, and WordNet can be visually inspected in Figure˜5 in Appendix˜G.
Table 1: Left: RSA comparisons of hierarchically sensitive pairwise similarities [52] in the unembedding spaces of VLM-LM pairs, and pairwise path-similarities from the WordNet (wn) Noun Hierarchy. Subscripts show standard deviation (hidden if under 0.01). Right: Differences ($\Delta$) in cosine similarities between positive concept pairs (i.e., in a hypernymy relationship) and negative samples from the taxonomy in TaxonomiGQA, computed using VLM and LM static-embedding layers.
@@ 4.2.2 Embedding similarities of taxonomic relations
Another way in which taxonomic relations can be investigated is via vector similarity—we tested whether the lexical embeddings (i.e., uncontextualized representations) corresponding to concepts in our reference taxonomy are more similar to embeddings of their hyponyms, relative to embeddings of non-hyponyms in our taxonomy. We computed the similarity between each target concept and its hyponym, as well as between the target concept and four randomly sampled non-hyponym concepts (same as in Section˜3.1). Then, we computed the difference between target-hyponym similarity and the average similarity between the target and the negative samples. We tested whether this difference is greater in VLMs than LMs, which would mean that VLM embeddings encode hypernym-hyponym relations more similarly than non-hypernym-hyponym relations. This hypothesis is not borne out: Table˜1 (right) shows that this holds for no VLM-LM pairs (the significant effect in Qwen2.5-I vs. Qwen2.5-VL-I is in the opposite direction).